# Measuring Social Norms of Large Language Models

### v2

@@ Appendix D Summary of Skills
Grade/Skill
Random
LLaMA2-70B-Chat
[…]
GPT3.5-Turbo
[…]
grade-8/is-the-pronoun-reflexive-or-intensive
43
64
70
57
92
grade-8/interpret-figures-of-speech
59
100
100
100
100
grade-8/identify-the-verb-mood
32
76
56
100
100
@@
grade-8/classify-figures-of-speech
64
100
84
100
78
grade-8/is-the-sentence-declarative-interrogative-imperative-or-exclamatory
36
83
94
100
100
grade-8/vocabulary-review-trace-an-argument
37
84
100
100
100
grade-8/is-it-a-direct-object-or-an-indirect-object
60
63
79
75
100
grade-8/identify-the-authors-purpose
34
100
100
100
100
grade-8/decide-whether-ellipses-are-used-appropriately
58
52
68
100
61
grade-8/compound-subjects-and-objects-with-pronouns
63
73
60
100
97
grade-8/identify-vague-pronoun-references
56
67
79
70
60
grade-8/choose-between-adjectives-and-adverbs
68
100
86
100
87
grade-8/what-does-the-punctuation-suggest
43
62
70
81
63
grade-8/suggest-appropriate-revisions
69
74
80
84
84
grade-8/identify-thesis-statements
79
65
75
75
82
grade-8/determine-the-themes-of-short-stories
36
59
80
59
69
grade-8/classify-logical-fallacies
55
67
52
100
100
grade-8/use-greek-and-latin-roots-as-clues-to-the-meanings-of-words
51
100
100
100
100
grade-8/interpret-the-meaning-of-an-allusion-from-its-source
63
100
100
100
100
grade-8/use-the-correct-frequently-confused-word
66
74
70
77
75
grade-8/positive-and-negative-connotation
67
70
91
100
100
grade-8/distinguish-facts-from-opinions
51
94
86
91
100
grade-8/use-thesaurus-entries
45
100
100
100
100
grade-8/use-the-correct-verb-with-compound-subjects
61
73
82
100
98
grade-8/determine-the-main-idea-of-a-passage
53
82
100
95
100
grade-8/analyze-short-stories
33
63
95
49
70
grade-8/identify-plurals-singular-possessives-and-plural-possessives
46
81
64
95
100
grade-8/determine-the-meanings-of-greek-and-latin-roots
28
84
79
100
100
grade-8/which-sentence-is-more-formal
64
100
100
100
100
grade-8/use-parallel-structure
55
64
69
92
94
grade-9/determine-the-meanings-of-words-with-greek-and-latin-roots
23
100
100
100
100
grade-9/use-words-as-clues-to-the-meanings-of-greek-and-latin-roots
25
100
73
100
100
grade-9/analogies
44
100
64
100
100
grade-9/is-it-a-phrase-or-a-clause
51
59
45
100
68
grade-9/is-the-sentence-simple-compound-complex-or-compound-complex
36
17
24
34
66
grade-9/analogies-challenge
37
78
82
100
100
grade-9/words-with-pre
54
100
100
100
100
grade-9/are-the-modifiers-used-correctly
38
80
84
82
85
grade-9/recall-the-source-of-an-allusion
49
100
100
100
100
grade-9/transitions-with-conjunctive-adverbs
55
80
67
91
77
grade-9/words-with-less
60
100
100
100
100
grade-9/identify-plagiarism
37
39
35
39
46
grade-9/words-with-re
25
100
100
100
100
grade-9/identify-the-narrative-point-of-view
36
52
49
82
82
grade-9/use-the-correct-pair-of-correlative-conjunctions
54
64
49
78
82
grade-9/identify-active-and-passive-voice
39
63
51
100
100
grade-9/interpret-figures-of-speech
62
100
100
100
100
grade-9/identify-the-verb-mood
36
75
66
100
100
grade-9/is-the-sentence-declarative-interrogative-imperative-or-exclamatory
13
73
85
100
79
grade-9/decide-whether-ellipses-are-used-appropriately
60
60
92
86
68
grade-9/identify-vague-pronoun-references
39
51
73
76
81
grade-9/choose-between-adjectives-and-adverbs
59
83
90
100
91
grade-9/what-does-the-punctuation-suggest
53
64
73
68
83
grade-9/suggest-appropriate-revisions
75
70
69
80
64
grade-9/identify-thesis-statements
48
77
46
64
73
grade-9/classify-logical-fallacies
46
95
50
100
94
grade-9/use-greek-and-latin-roots-as-clues-to-the-meanings-of-words
41
100
100
100
100
grade-9/interpret-the-meaning-of-an-allusion-from-its-source
57
100
100
100
100
grade-9/use-the-correct-frequently-confused-word
52
68
74
88
75
grade-9/words-with-mis
62
95
100
100
100
grade-9/distinguish-facts-from-opinions
66
83
100
100
100
grade-9/use-thesaurus-entries
40
100
100
100
100
grade-9/determine-the-main-idea-of-a-passage
79
100
100
100
100
grade-9/determine-the-meanings-of-greek-and-latin-roots
28
82
76
100
100
grade-9/use-parallel-structure
69
87
65
100
77
grade-9/use-dictionary-entries-to-determine-correct-usage
72
85
100
72
93
grade-9/which-text-is-most-formal
58
100
100
100
100
grade-9/choose-the-topic-sentence-that-best-captures-the-main-idea
58
38
100
97
100
grade-9/identify-sentence-fragments
46
54
41
100
90
grade-9/use-context-as-a-clue-to-the-meanings-of-foreign-expressions
36
100
100
100
100
grade-9/use-the-correct-foreign-expression
69
100
100
100
100
grade-9/identify-audience-and-purpose
46
100
100
100
100
grade-9/classify-figures-of-speech-euphemism-hyperbole-oxymoron-paradox
65
65
84
84
92
grade-9/explore-words-with-new-or-contested-usages
69
57
64
60
54
grade-9/analyze-short-stories-set-1
32
93
84
59
100
grade-9/choose-punctuation-to-avoid-fragments-and-run-ons
18
22
39
49
53
grade-9/transition-logically-between-claims-evidence-analysis-and-counterclaims
41
94
79
91
83
grade-9/choose-the-analysis-that-logically-connects-the-evidence-to-the-claim
54
88
100
100
86
grade-9/identify-run-on-sentences
79
63
63
80
73
grade-9/use-words-accurately-and-precisely
59
85
65
100
92
grade-9/classify-figures-of-speech-review
43
100
95
100
83
grade-9/avoid-double-illogical-and-unclear-comparisons
37
46
66
69
88
grade-9/choose-the-best-evidence-to-support-a-claim
19
24
67
37
100
grade-9/use-appeals-to-ethos-pathos-and-logos-in-persuasive-writing
49
50
66
35
67
grade-9/identify-elements-of-poetry
54
56
70
76
87
grade-9/choose-the-word-whose-connotation-and-denotation-best-match-the-sentence
25
100
100
100
100
grade-9/analyze-short-stories-set-2
55
44
84
66
100
grade-9/evaluate-counterclaims
60
56
100
60
100
grade-9/use-etymologies-to-determine-the-meanings-of-words
41
100
100
100
100
grade-10/determine-the-meanings-of-words-with-greek-and-latin-roots
33
100
100
100
100
grade-10/use-words-as-clues-to-the-meanings-of-greek-and-latin-roots
36
100
90
100
100
grade-10/analogies
29
67
67
88
100
grade-10/is-it-a-phrase-or-a-clause
62
67
72
89
79
grade-10/is-the-sentence-simple-compound-complex-or-compound-complex
38
23
26
41
75
grade-10/analogies-challenge
39
85
45
100
79
grade-10/are-the-modifiers-used-correctly
55
73
60
95
81
grade-10/recall-the-source-of-an-allusion
48
95
100
100
100
grade-10/transitions-with-conjunctive-adverbs
41
79
64
70
83
grade-10/identify-plagiarism
23
13
14
23
38
grade-10/use-the-correct-pair-of-correlative-conjunctions
64
70
59
76
88
grade-10/identify-active-and-passive-voice
68
79
43
100
100
grade-10/interpret-figures-of-speech
35
100
100
100
100
grade-10/is-the-sentence-declarative-interrogative-imperative-or-exclamatory
26
49
80
100
84
grade-10/decide-whether-ellipses-are-used-appropriately
64
42
80
80
62
grade-10/identify-vague-pronoun-references
77
64
81
62
71
grade-10/choose-between-adjectives-and-adverbs
36
84
82
97
82
grade-10/what-does-the-punctuation-suggest
39
62
73
61
83
grade-10/suggest-appropriate-revisions
52
64
73
68
73
grade-10/identify-thesis-statements
65
33
34
71
67
grade-10/classify-logical-fallacies
36
100
60
100
100
grade-10/use-greek-and-latin-roots-as-clues-to-the-meanings-of-words
40
100
100
100
100
grade-10/interpret-the-meaning-of-an-allusion-from-its-source
60
100
100
100
100
grade-10/use-the-correct-frequently-confused-word
58
63
56
100
75
@@
Grade/Skill
Random
LLaMA2-70B-Chat
[…]
GPT3.5-Turbo
[…]
grade-10/use-thesaurus-entries
68
100
100
100
100
grade-10/determine-the-main-idea-of-a-passage
53
100
11
100
100
grade-10/determine-the-meanings-of-greek-and-latin-roots
28
100
90
100
100
grade-10/use-parallel-structure
76
75
66
70
100
grade-10/use-dictionary-entries-to-determine-correct-usage
55
18
85
49
93
grade-10/which-text-is-most-formal
40
100
100
100
100
@@
grade-10/choose-the-topic-sentence-that-best-captures-the-main-idea
37
39
93
100
100
grade-10/identify-sentence-fragments
42
45
42
100
95
grade-10/use-context-as-a-clue-to-the-meanings-of-foreign-expressions
26
100
100
100
100
grade-10/use-the-correct-foreign-expression
73
100
100
100
100
grade-10/identify-audience-and-purpose
40
100
71
90
90
@@
grade-10/classify-figures-of-speech-euphemism-hyperbole-oxymoron-paradox
63
89
80
91
83
grade-10/explore-words-with-new-or-contested-usages
49
65
50
59
63
grade-10/choose-punctuation-to-avoid-fragments-and-run-ons
37
41
41
34
36
@@
grade-10/transition-logically-between-claims-evidence-analysis-and-counterclaims
45
74
64
100
90
@@
grade-10/choose-the-analysis-that-logically-connects-the-evidence-to-the-claim
37
87
100
100
100
grade-10/identify-run-on-sentences
48
64
53
100
92
grade-10/avoid-double-illogical-and-unclear-comparisons
15
64
71
67
79
grade-10/use-words-accurately-and-precisely
43
86
73
100
86
grade-10/classify-figures-of-speech-review
44
91
100
85
66
grade-10/choose-the-best-evidence-to-support-a-claim
61
21
100
68
100
grade-10/use-appeals-to-ethos-pathos-and-logos-in-persuasive-writing
37
64
52
37
84
@@
grade-10/choose-the-word-whose-connotation-and-denotation-best-match-the-sentence
63
100
100
100
100
grade-10/use-etymologies-to-determine-the-meanings-of-words
42
100
100
100
100
@@
grade-11/determine-the-meanings-of-words-with-greek-and-latin-roots
36
100
100
100
100
grade-11/describe-the-difference-between-related-words
46
100
100
100
100

### v3

@@ Appendix D Summary of Skills
Grade/Skill Random LLaMA2-70B-Chat SocialAgentLLaMA2-70B-Chat GPT3.5-Turbo SocialAgentGPT3.5-Turbo
grade-8/is-the-pronoun-reflexive-or-intensive 43 64 70 57 92
grade-8/interpret-figures-of-speech 59 100 100 100 100
grade-8/identify-the-verb-mood 32 76 56 100 100
grade-8/classify-figures-of-speech 64 100 84 100 78
grade-8/is-the-sentence-declarative-interrogative-imperative-or-exclamatory 36 83 94 100 100
grade-8/vocabulary-review-trace-an-argument 37 84 100 100 100
grade-8/is-it-a-direct-object-or-an-indirect-object 60 63 79 75 100
grade-8/identify-the-authors-purpose 34 100 100 100 100
grade-8/decide-whether-ellipses-are-used-appropriately 58 52 68 100 61
grade-8/compound-subjects-and-objects-with-pronouns 63 73 60 100 97
grade-8/identify-vague-pronoun-references 56 67 79 70 60
grade-8/choose-between-adjectives-and-adverbs 68 100 86 100 87
grade-8/what-does-the-punctuation-suggest 43 62 70 81 63
grade-8/suggest-appropriate-revisions 69 74 80 84 84
grade-8/identify-thesis-statements 79 65 75 75 82
grade-8/determine-the-themes-of-short-stories 36 59 80 59 69
grade-8/classify-logical-fallacies 55 67 52 100 100
grade-8/use-greek-and-latin-roots-as-clues-to-the-meanings-of-words 51 100 100 100 100
grade-8/interpret-the-meaning-of-an-allusion-from-its-source 63 100 100 100 100
grade-8/use-the-correct-frequently-confused-word 66 74 70 77 75
grade-8/positive-and-negative-connotation 67 70 91 100 100
grade-8/distinguish-facts-from-opinions 51 94 86 91 100
grade-8/use-thesaurus-entries 45 100 100 100 100
grade-8/use-the-correct-verb-with-compound-subjects 61 73 82 100 98
grade-8/determine-the-main-idea-of-a-passage 53 82 100 95 100
grade-8/analyze-short-stories 33 63 95 49 70
grade-8/identify-plurals-singular-possessives-and-plural-possessives 46 81 64 95 100
grade-8/determine-the-meanings-of-greek-and-latin-roots 28 84 79 100 100
grade-8/which-sentence-is-more-formal 64 100 100 100 100
grade-8/use-parallel-structure 55 64 69 92 94
grade-9/determine-the-meanings-of-words-with-greek-and-latin-roots 23 100 100 100 100
grade-9/use-words-as-clues-to-the-meanings-of-greek-and-latin-roots 25 100 73 100 100
grade-9/analogies 44 100 64 100 100
grade-9/is-it-a-phrase-or-a-clause 51 59 45 100 68
grade-9/is-the-sentence-simple-compound-complex-or-compound-complex 36 17 24 34 66
grade-9/analogies-challenge 37 78 82 100 100
grade-9/words-with-pre 54 100 100 100 100
grade-9/are-the-modifiers-used-correctly 38 80 84 82 85
grade-9/recall-the-source-of-an-allusion 49 100 100 100 100
grade-9/transitions-with-conjunctive-adverbs 55 80 67 91 77
grade-9/words-with-less 60 100 100 100 100
grade-9/identify-plagiarism 37 39 35 39 46
grade-9/words-with-re 25 100 100 100 100
grade-9/identify-the-narrative-point-of-view 36 52 49 82 82
grade-9/use-the-correct-pair-of-correlative-conjunctions 54 64 49 78 82
grade-9/identify-active-and-passive-voice 39 63 51 100 100
grade-9/interpret-figures-of-speech 62 100 100 100 100
grade-9/identify-the-verb-mood 36 75 66 100 100
grade-9/is-the-sentence-declarative-interrogative-imperative-or-exclamatory 13 73 85 100 79
grade-9/decide-whether-ellipses-are-used-appropriately 60 60 92 86 68
grade-9/identify-vague-pronoun-references 39 51 73 76 81
grade-9/choose-between-adjectives-and-adverbs 59 83 90 100 91
grade-9/what-does-the-punctuation-suggest 53 64 73 68 83
grade-9/suggest-appropriate-revisions 75 70 69 80 64
grade-9/identify-thesis-statements 48 77 46 64 73
grade-9/classify-logical-fallacies 46 95 50 100 94
grade-9/use-greek-and-latin-roots-as-clues-to-the-meanings-of-words 41 100 100 100 100
grade-9/interpret-the-meaning-of-an-allusion-from-its-source 57 100 100 100 100
grade-9/use-the-correct-frequently-confused-word 52 68 74 88 75
grade-9/words-with-mis 62 95 100 100 100
grade-9/distinguish-facts-from-opinions 66 83 100 100 100
grade-9/use-thesaurus-entries 40 100 100 100 100
grade-9/determine-the-main-idea-of-a-passage 79 100 100 100 100
grade-9/determine-the-meanings-of-greek-and-latin-roots 28 82 76 100 100
grade-9/use-parallel-structure 69 87 65 100 77
grade-9/use-dictionary-entries-to-determine-correct-usage 72 85 100 72 93
grade-9/which-text-is-most-formal 58 100 100 100 100
grade-9/choose-the-topic-sentence-that-best-captures-the-main-idea 58 38 100 97 100
grade-9/identify-sentence-fragments 46 54 41 100 90
grade-9/use-context-as-a-clue-to-the-meanings-of-foreign-expressions 36 100 100 100 100
grade-9/use-the-correct-foreign-expression 69 100 100 100 100
grade-9/identify-audience-and-purpose 46 100 100 100 100
@@
grade-9/classify-figures-of-speech-euphemism-hyperbole-oxymoron-paradox 65 65 84 84 92
grade-9/explore-words-with-new-or-contested-usages 69 57 64 60 54
grade-9/analyze-short-stories-set-1 32 93 84 59 100
grade-9/choose-punctuation-to-avoid-fragments-and-run-ons 18 22 39 49 53
grade-9/transition-logically-between-claims-evidence-analysis-and-counterclaims 41 94 79 91 83
grade-9/choose-the-analysis-that-logically-connects-the-evidence-to-the-claim 54 88 100 100 86
grade-9/identify-run-on-sentences 79 63 63 80 73
grade-9/use-words-accurately-and-precisely 59 85 65 100 92
grade-9/classify-figures-of-speech-review 43 100 95 100 83
grade-9/avoid-double-illogical-and-unclear-comparisons 37 46 66 69 88
grade-9/choose-the-best-evidence-to-support-a-claim 19 24 67 37 100
grade-9/use-appeals-to-ethos-pathos-and-logos-in-persuasive-writing 49 50 66 35 67
grade-9/identify-elements-of-poetry 54 56 70 76 87
grade-9/choose-the-word-whose-connotation-and-denotation-best-match-the-sentence 25 100 100 100 100
grade-9/analyze-short-stories-set-2 55 44 84 66 100
grade-9/evaluate-counterclaims 60 56 100 60 100
grade-9/use-etymologies-to-determine-the-meanings-of-words 41 100 100 100 100
grade-10/determine-the-meanings-of-words-with-greek-and-latin-roots 33 100 100 100 100
grade-10/use-words-as-clues-to-the-meanings-of-greek-and-latin-roots 36 100 90 100 100
grade-10/analogies 29 67 67 88 100
grade-10/is-it-a-phrase-or-a-clause 62 67 72 89 79
grade-10/is-the-sentence-simple-compound-complex-or-compound-complex 38 23 26 41 75
grade-10/analogies-challenge 39 85 45 100 79
grade-10/are-the-modifiers-used-correctly 55 73 60 95 81
grade-10/recall-the-source-of-an-allusion 48 95 100 100 100
grade-10/transitions-with-conjunctive-adverbs 41 79 64 70 83
grade-10/identify-plagiarism 23 13 14 23 38
grade-10/use-the-correct-pair-of-correlative-conjunctions 64 70 59 76 88
grade-10/identify-active-and-passive-voice 68 79 43 100 100
grade-10/interpret-figures-of-speech 35 100 100 100 100
grade-10/is-the-sentence-declarative-interrogative-imperative-or-exclamatory 26 49 80 100 84
grade-10/decide-whether-ellipses-are-used-appropriately 64 42 80 80 62
grade-10/identify-vague-pronoun-references 77 64 81 62 71
grade-10/choose-between-adjectives-and-adverbs 36 84 82 97 82
grade-10/what-does-the-punctuation-suggest 39 62 73 61 83
grade-10/suggest-appropriate-revisions 52 64 73 68 73
grade-10/identify-thesis-statements 65 33 34 71 67
grade-10/classify-logical-fallacies 36 100 60 100 100
grade-10/use-greek-and-latin-roots-as-clues-to-the-meanings-of-words 40 100 100 100 100
grade-10/interpret-the-meaning-of-an-allusion-from-its-source 60 100 100 100 100
grade-10/use-the-correct-frequently-confused-word 58 63 56 100 75
@@
Grade/Skill Random LLaMA2-70B-Chat SocialAgentLLaMA2-70B-Chat GPT3.5-Turbo SocialAgentGPT3.5-Turbo
grade-10/use-thesaurus-entries 68 100 100 100 100
grade-10/determine-the-main-idea-of-a-passage 53 100 11 100 100
grade-10/determine-the-meanings-of-greek-and-latin-roots 28 100 90 100 100
grade-10/use-parallel-structure 76 75 66 70 100
grade-10/use-dictionary-entries-to-determine-correct-usage 55 18 85 49 93
grade-10/which-text-is-most-formal 40 100 100 100 100
@@
grade-10/choose-the-topic-sentence-that-best-captures-the-main-idea 37 39 93 100 100
grade-10/identify-sentence-fragments 42 45 42 100 95
grade-10/use-context-as-a-clue-to-the-meanings-of-foreign-expressions 26 100 100 100 100
grade-10/use-the-correct-foreign-expression 73 100 100 100 100
grade-10/identify-audience-and-purpose 40 100 71 90 90
@@
grade-10/classify-figures-of-speech-euphemism-hyperbole-oxymoron-paradox 63 89 80 91 83
grade-10/explore-words-with-new-or-contested-usages 49 65 50 59 63
grade-10/choose-punctuation-to-avoid-fragments-and-run-ons 37 41 41 34 36
@@
grade-10/transition-logically-between-claims-evidence-analysis-and-counterclaims 45 74 64 100 90
@@
grade-10/choose-the-analysis-that-logically-connects-the-evidence-to-the-claim 37 87 100 100 100
grade-10/identify-run-on-sentences 48 64 53 100 92
grade-10/avoid-double-illogical-and-unclear-comparisons 15 64 71 67 79
grade-10/use-words-accurately-and-precisely 43 86 73 100 86
grade-10/classify-figures-of-speech-review 44 91 100 85 66
grade-10/choose-the-best-evidence-to-support-a-claim 61 21 100 68 100
grade-10/use-appeals-to-ethos-pathos-and-logos-in-persuasive-writing 37 64 52 37 84
@@
grade-10/choose-the-word-whose-connotation-and-denotation-best-match-the-sentence 63 100 100 100 100
grade-10/use-etymologies-to-determine-the-meanings-of-words 42 100 100 100 100
@@
grade-11/determine-the-meanings-of-words-with-greek-and-latin-roots 36 100 100 100 100
grade-11/describe-the-difference-between-related-words 46 100 100 100 100
grade-11/use-words-as-clues-to-the-meanings-of-greek-and-latin-roots 35 100 100 100 100
grade-11/analogies 54 69 67 92 89
grade-11/is-it-a-phrase-or-a-clause 45 45 45 78 76
grade-11/is-the-sentence-simple-compound-complex-or-compound-complex 27 24 34 63 43
grade-11/analogies-challenge 36 93 59 88 77
grade-11/are-the-modifiers-used-correctly 39 60 58 72 86
grade-11/recall-the-source-of-an-allusion 55 100 100 100 100
grade-11/transitions-with-conjunctive-adverbs 50 64 77 87 100
grade-11/decide-whether-ellipses-are-used-appropriately 78 62 77 90 71
grade-11/identify-vague-pronoun-references 49 58 64 62 87
grade-11/choose-between-adjectives-and-adverbs 59 95 69 80 94
grade-11/what-does-the-punctuation-suggest 59 50 60 73 76
grade-11/suggest-appropriate-revisions 51 58 59 56 64
grade-11/classify-logical-fallacies 39 67 55 100 88
grade-11/use-greek-and-latin-roots-as-clues-to-the-meanings-of-words 40 100 100 100 100
grade-11/interpret-the-meaning-of-an-allusion-from-its-source 60 100 100 100 100
grade-11/use-the-correct-frequently-confused-word 64 64 59 100 88
grade-11/determine-the-main-idea-of-a-passage 47 100 100 100 100
grade-11/determine-the-meanings-of-greek-and-latin-roots 33 100 79 100 100
grade-11/use-parallel-structure 100 100 61 79 100
grade-11/use-dictionary-entries-to-determine-correct-usage 67 51 100 72 100
grade-11/which-text-is-most-formal 69 97 100 100 100
grade-11/identify-sentence-fragments 75 62 49 100 100
grade-11/use-context-as-a-clue-to-the-meanings-of-foreign-expressions 30 100 100 100 100
grade-11/use-the-correct-foreign-expression 77 100 100 100 100
grade-11/identify-audience-and-purpose 14 100 99 100 100
grade-11/explore-words-with-new-or-contested-usages 73 50 55 42 67
grade-11/analyze-short-stories-set-1 51 62 92 100 77
grade-11/choose-punctuation-to-avoid-fragments-and-run-ons 28 35 40 73 60
grade-11/transition-logically-between-claims-evidence-analysis-and-counterclaims 40 87 54 83 69
grade-11/choose-the-analysis-that-logically-connects-the-evidence-to-the-claim 44 90 90 100 86
grade-11/identify-run-on-sentences 40 65 59 90 86
grade-11/use-words-accurately-and-precisely 44 100 86 100 100
grade-11/avoid-double-illogical-and-unclear-comparisons 44 64 68 63 82
grade-11/choose-the-best-evidence-to-support-a-claim 29 15 76 71 100
grade-11/use-appeals-to-ethos-pathos-and-logos-in-persuasive-writing 48 52 45 58 63
grade-11/choose-the-word-whose-connotation-and-denotation-best-match-the-sentence 55 100 100 100 100
grade-11/evaluate-counterclaims 33 58 100 36 100
grade-11/use-etymologies-to-determine-the-meanings-of-words 62 100 100 100 100
grade-11/classify-the-figure-of-speech-euphemism-hyperbole-oxymoron-paradox 48 82 67 100 88
grade-11/classify-the-figure-of-speech-anaphora-antithesis-apostrophe-assonance-chiasmus-understatement 64 100 100 100 100
grade-11/interpret-the-figure-of-speech 45 100 100 100 100
grade-11/prefixes 61 100 84 100 100
grade-11/suffixes 58 100 100 100 100
grade-11/classify-the-figure-of-speech-review 71 100 79 100 100